# A neural network-based approach to hybrid systems identification for control

### v1

@@ Abstract
We consider the problem of designing a machine learning-based model of an unknown dynamical system from a finite number of (state-input)-successor state data points, such that the model obtained is also suitable for optimal control design. We propose a specific neural network (NN) architecture that yields a hybrid system with piecewise-affine (PWA) dynamics that is differentiable with respect to the network’s parameters, thereby enabling the use of derivative-based training procedures. We show that a careful choice of our NN’s weights produces a hybrid system model with structural properties that are highly favourable when used as part of a finite horizon optimal control problem (OCP). Specifically, we show that optimal solutions with strong local optimality guarantees can be computed via nonlinear programming (NLP), in contrast to classical OCPs for general hybrid systems which typically require mixed-integer optimization. In addition to being well-suited for optimal control design, numerical simulations illustrate that our NN-based technique enjoys very similar performance to state-of-the-art system identification methodologies for hybrid systems and it is competitive on nonlinear benchmarks.
thanks: Email addresses: {filippo.fabiani, daniele.masti}@imtlucca.it, bstellato@princeton.edu, paul.goulart@eng.ox.ac.uk.
@@ 1 Introduction
Interest in data-driven learning techniques is reshaping conventional systems modelling and controller design, with the convergence of accessible datasets and advanced machine learning methods offering a transformative opportunity [4].
Within the systems-and-control community, systems identification methods have benefited significantly from powerful tools drawn from the machine learning literature [55, 13, 58, 6]. In particular, NNs [30] have been used since at least the 1980s [62] to generate data-based system model surrogates, which has sparked renewed interest in black-box modelling of functions within the state–space. Their flexibility, given both their universal approximation properties [41] and the many NNs architectures available, has motivated research towards the development of NN-based techniques for (non)linear system identification adopting various neural model structures [24, 22, 23, 3, 52, 51].
However, with a few exceptions [40, 48] available learning-based system identification techniques do not consider whether the models constructed will have an internal structure favourable for use within an optimal control problem. In fact, unless one assumes a certain simplified structure, deep NNs are generally hard to analyze due to their nonlinear and large-scale structure [31, 30]. As an undesirable consequence, adopting the resulting system model surrogates into OCPs may make the latter extremely hard to solve efficiently (wherever a solution is guaranteed to exist). In this paper we propose a NN-based technique to identify a hybrid system representation in the form of continuous PWA dynamics with structure tailored specifically for optimal control design.
@@ 1.1 Related work
Although hybrid models represent a broad class of systems [54, 29] whose identification is known to be NP-hard in general [57, 46], to the best of our knowledge little attention has focussed on the derivation of hybrid system models through (deep) NN-based surrogates.
Starting from the pioneering work in [5], where a continuous PWA representation was obtained through a NN with a specific two-layer structure on the basis of the results developed in [16], [15] constructed a PWA local model by introducing an input space tessellation via hyperpolyhedral convex cells, associating to them NN granules with a local interpolation capability. The authors of [42] instead introduced a constructive algorithm that builds a NN with PWA sigmoidal nodes and one hidden layer. In particular, one node at a time was added by fitting the residual, a task that was accomplished by searching for the best fit through a sequence of quadratic programs (QPs). A single hidden layer network was also proposed in [27], which constructed a continuous PWA error function and developed an efficient, two-step algorithm working in the weight space to minimize it. More recently, [21] proposed a series of experiments in which newly developed libraries are employed to identify dynamical models with NNs even for complicated hybrid systems, while more applied contributions have analyzed the positive effects of imposing structured knowledge to deep NN-based surrogates in describing the multi-modal behaviour exhibited by the system at hand [18, 47].
In a similar spirit to our work, [44] proposed to infer a control-oriented linear complementarity (LC) dynamics (equivalent to a PWA one [35]) from data directly. In particular, their approach built upon the minimization of a tailored violation-based loss function, which allowed one to learn LC dynamics via standard gradient-based methods. In view of the results established later in the present paper, we also mention the study conducted in [63]. There, a (strong) stationarity condition was identified as necessary and sufficient to characterize local optima of a program with a rectified linear unit (ReLU) NN entering both in the cost and constraints, and the corresponding LC-based, lifted reformulation yielding an mathematical program with complementarity constraints (MPCC).
@@ 1.2 Summary of contribution
Following from our recent work [19, 20], in this paper we take a NN-based approach to the problem of system identification for control. In particular, we wish to answer the following question: given $N$ (state-input)-successor state samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, how can we obtain a descriptive system model that is also suitable for optimal control design? While the meaning of the terms “descriptive” and “suitable” will be clarified later in the paper, we summarize the contribution as follows:
@@
    We employ a NN-based method to obtain a hybrid system with continuous PWA dynamics from available data. In particular, the proposed NN is characterized by a simple architecture, and its output is differentiable with respect to (w.r.t.) the NN’s parameters;
@@
    Given its end-to-end trainability, a careful choice of some weights of the NN allows us to produce a hybrid dynamics model with a specific structure. We then show that the latter can be controlled (locally) optimally by solving the Karush-Kuhn-Tucker (KKT) conditions of the underlying finite horizon OCP;
@@
Our method thus requires a simple identification step, represented by a careful training of a NN with specific structure via standard tools, which yields a hybrid system model that is well-suited to optimal control design. In particular, solving an OCP involving a hybrid system with PWA dynamics as an NLP has been proven to require shorter computation times and feature better scaling in the problem dimensions than standard approaches based on mixed-integer optimization [38, 32]. In addition, we recall that PWA regression is NP-hard in general [46], since it requires simultaneous classification of the $N$ samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ into modes, thereby calling for a regression of a submodel for each mode. By taking a NN-based perspective, we instead circumvent any challenge resulting from (possibly expensive) identification procedures, since the output differentiability of the proposed NN architecture allows for a standard training procedure requiring little computational efforts.
Unlike other approaches, we obtain an LC model suitable for control as a direct consequence of a careful choice of a NN with specific architecture, thereby circumventing the requirement of strict complementarity to recover differentiability w.r.t. the main parameters as in [44, 18]. In contrast to [63], we provide an explicit structure of a NN for which local stationarity conditions, coinciding with the standard KKT system, are known to hold for the MPCC obtained by using our particular NN as a hybrid system model in an OCP.
The rest of the paper is organized as follows: we formalize the problem addressed in §2, and in §3 introduce our NN-based approach to system identification for control. In §4 we describe the OCP obtained when using our identified model structure, and develop our main theoretical results relating to that problem in §5. Finally, §6 reports a series of numerical experiments.
@@ 2 Problem formulation
We will assume that we have available a finite collection of (state-input)-successor state measured triplets, $(x,u,x^{+})$, $x$, $x^{+}\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$ for an unknown but deterministic dynamic system. Our first aim is to produce a model of this unknown system without running further experiments or successive identification procedures. In particular, the model we wish to compute from the available data should be “descriptive enough”, i.e., it should belong to a model class capable of capturing a wide variety of system behaviours, while at the same time being suited for optimal control.
To clarify what we mean with the latter statement, we introduce next the finite horizon optimal control problem (OCP) considered throughout the paper:
@@
where $x_{t}$, $t\in\mathcal{T}\coloneqq\{0,\ldots,T-1\}$, denotes the predicted system state $t$ timesteps into the future through the (possibly nonlinear) dynamical model $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, which is to be inferred from data and is characterized by parameters $\theta\in\mathbb{R}^{p}$. We denote by $\boldsymbol{u}\coloneqq\textrm{col}((u_{t})_{t\in\mathcal{T}})\in\mathbb{R}^{mT}$ the collection of control inputs to be chosen and by $\boldsymbol{x}\coloneqq\textrm{col}((x_{t})_{t\in\mathcal{T}\cup\{T\}})\in\mathbb{R}^{n(T+1)}$ the resulting state trajectory starting from the measured initial state $x_{0}=x(0)$. In addition, let $\ell_{t}$ and $\ell_{T}$ denote the stage and terminal costs respectively, which for technical reasons are assumed to meet the following standard condition:
@@
Therefore, given $N$ samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, our goal is to develop a system model $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that, once $\mathcal{N}_{\theta}$ is used as part of the OCP in (1), an optimal solution to the latter can be characterized via standard optimality conditions. To this end, we choose for $\mathcal{N}_{\theta}$ a neural network (NN)-based representation of a hybrid system, ultimately leading to an OCP equivalent to (1) in the form of a mathematical program with complementarity constraints (MPCC) [49], whose classical Karush-Kuhn-Tucker (KKT) conditions [8, §5.1] will be necessary and sufficient to characterize an optimal solution.
For technical reasons, we will assume in the remainder that the state-input pair takes values in some bounded set as follows:
@@
3 Neural network-based representation of hybrid systems
@@
To address the problem introduced in §2, we will design $\mathcal{N}_{\theta}$ as a two-layer NN with the simplified architecture illustrated in Fig. 1, and then make use of available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ to train the associated parameters, generically described by $\theta\in\mathbb{R}^{p}$. In particular, the underlying NN consists of an OptNet layer [2, 1], which takes the input pair $\textrm{col}(x,u)\eqqcolon y$ as a parameter to solve the following quadratic program (QP):
@@
thus returning an optimal solution $z^{\star}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{s}$ as output. This output is then passed through a buffer layer performing an affine transformation so that
$x^{+}=Wz^{\star}(x,u)+c\eqqcolon\mathcal{N}_{\theta}(x,u),$ (3)
with appropriate matrix $W\in\mathbb{R}^{n\times s}$ and affine term $c\in\mathbb{R}^{n}$. The parameters characterizing $\mathcal{N}_{\theta}$ are then $\theta\coloneqq\{(Q,q,A,b,G,h),W,c\}\in\mathbb{R}^{p}$, which must be determined during some offline training procedure using our available data set $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$.
To ensure a unique optimizer for (2) so that there is no ambiguity in the definition of the dynamics in (3), during the training phase we will impose strict convexity of (2), i.e., $Q\in\mathbb{S}^{s}_{\succ 0}$. This will be made possible in view of the main technical features possessed by the NN $\mathcal{N}_{\theta}$ illustrated in Fig. 1, which are established next:
@@
The NN $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ in (3), consisting of an OptNet layer and a affine buffer layer in cascade, enjoys the following properties:
@@
(ii) This part follows from Standing Assumption 2 and [37, Th. 2] directly. In fact, if one chooses $c=0$, and parameters $Q,q,A,b,G,h$ so that (2) becomes:
@@
with $Q\in\mathbb{S}^{s}_{\succ 0}$, then the structure in [37, Eq. 4] is immediately recovered. Note that these choices are always possible since differentiability of the NN output w.r.t. $\theta$, which follows from the first property shown in this statement, implies that $\mathcal{N}_{\theta}$ is an end-to-end trainable NN. ∎
Proposition 1(i) thus makes $\mathcal{N}_{\theta}$ in (3), with $z^{\star}(\cdot)$ solving (2), an end-to-end trainable NN, meaning that one is actually able to set values for the main parameters $\theta=\{(Q,q,A,b,G,h),W,c\}$ exploiting (state-input)-successor state samples, as commonly happens with any other NN. A differentiable output enables the use of a typical backpropagation strategy to find the gradient of the training loss function. Consequently, one can train $\mathcal{N}_{\theta}$ with standard gradient descent methods for finding a (local) minimum of that function.
As a main consequence of the second property in Proposition 1, instead, we note that (1) turns out to coincide with an OCP involving a continuous PWA dynamics. The latter is known to be equivalent to a number of hybrid system model classes [35], such as linear complementarity (LC) [34, 44] or mixed-logical dynamical (MLD) models [7]. For this reason, the family of continuous PWA dynamics excels in capturing a wide variety of real-world system behaviours. In particular, since $\mathcal{N}_{\theta}$ will be trained on the basis of available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, (1) actually amounts to an OCP involving a data-based hybrid system representation.

### v2

@@ Abstract
We consider the problem of designing a machine learning-based model of an unknown dynamical system from a finite number of (state-input)-successor state data points, such that the model obtained is also suitable for optimal control design. We adopt a neural network (NN) architecture that, once suitably trained, yields a hybrid system with continuous piecewise-affine (PWA) dynamics that is differentiable with respect to the network’s parameters, thereby enabling the use of derivative-based training procedures. We show that a careful choice of our NN’s weights produces a hybrid system model with structural properties that are highly favorable when used as part of a finite horizon optimal control problem (OCP). Specifically, we rely on available results to establish that optimal solutions with strong local optimality guarantees can be computed via nonlinear programming (NLP), in contrast to classical OCPs for general hybrid systems which typically require mixed-integer optimization. Besides being well-suited for optimal control design, numerical simulations illustrate that our NN-based technique enjoys very similar performance to state-of-the-art system identification methods for hybrid systems and it is competitive on nonlinear benchmarks.
thanks: Email addresses: filippo.fabiani@imtlucca.it, bstellato@princeton.edu, daniele.masti@gssi.it, paul.goulart@eng.ox.ac.uk.
@@ 1 Introduction
Within the systems-and-control community, system identification methods have greatly benefited from powerful tools originating in the machine learning literature [49, 11, 5]. In particular, NNs [27] have been used since the 1980s [54] to produce data-driven system model surrogates. Their flexibility—stemming from their universal approximation capability [36] and the wide range of available architectures—has motivated research into NN-based techniques for (non)linear system identification adopting various structures [22, 21, 3, 46, 45].
Despite a few exceptions [35, 43], available learning-based system identification techniques do not consider whether the resulting models possess an internal structure favourable for use within an OCP. In fact, unless one assumes a certain simplified structure, deep NNs are generally hard to analyze due to their nonlinear and large-scale structure [27]. As an undesirable consequence, adopting the resulting system model surrogates into OCPs may make the latter extremely hard to solve efficiently (wherever a solution is guaranteed to exist). In this paper we propose a NN-based technique to identify a hybrid system representation, in the form of continuous PWA dynamics, with specific structure suitable for optimal control design.
@@ 1.1 Related work
Although hybrid models represent a broad class of systems [26] whose identification is known to be NP-hard in general [51, 41], to the best of our knowledge little attention has focused on the derivation of such models through (deep) NN-based surrogates.
Starting from the pioneering work in [4], where a continuous PWA representation was obtained through a NN with a specific two-layer structure by relying on [14], [13] constructed a PWA local model by introducing an input space tessellation via hyperpolyhedral convex cells, associating to them NN granules with a local interpolation capability. In [37], instead, the authors introduced a constructive algorithm that builds a NN with PWA sigmoidal nodes and one hidden layer. In particular, one node at a time was added by fitting the residual, a task that was accomplished by searching for the best fit through a sequence of quadratic programs (QPs). A single hidden layer network was also proposed in [24], which constructed a continuous PWA error function and developed an efficient, two-step algorithm working in the weight space to minimize it. Recently, [19] proposed a series of experiments in which novel libraries are employed to identify dynamical models with NNs for complicated hybrid systems, while [16, 42] have analyzed the effects of exploiting structured knowledge to NN surrogates in describing the system multi-modal behaviour.
Similar to our approach, [39] proposed to infer a control-oriented linear complementarity (LC) dynamics (equivalent to a PWA one [30]) directly from data. In particular, the approach in [39] built upon the minimization of a tailored violation-based loss function, which allowed one to learn LC dynamics via standard gradient-based methods. In view of our results, we also mention the approach in [55]. There, a stationarity condition was identified as necessary and sufficient to characterize local optima of a program with a rectified linear unit (ReLU) NN entering both in the cost and constraints, and the corresponding LC-based reformulation yielding an mathematical program with complementarity constraints (MPCC).
@@ 1.2 Summary of contribution
Following our recent contributions [17, 18], in this paper we take a NN-based approach to the problem of system identification for control. In particular, we wish to answer the following question: given $N$ (state-input)-successor state samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, how can we obtain a descriptive system model that is also suitable for optimal control design? While the meaning of “descriptive” and “suitable” will be clarified later in the paper, we summarize the contribution as follows:
@@
    We employ a NN-based method to obtain a hybrid system with continuous PWA dynamics from available data. The simple NN architecture we adopt combines an OptNet layer [2, 1] and an affine one, and results in a differentiable output with respect to (w.r.t.) the NN’s parameters;
@@
    Given its end-to-end trainability, we show that a careful choice of some weights of the NN allows us to produce a hybrid dynamics model with a specific structure. By relying on the results in [33], we then establish that the latter can be controlled (locally) optimally by solving the Karush-Kuhn-Tucker (KKT) conditions of the underlying OCP;
@@
Our method thus requires a simple identification step, represented by a careful training of a NN with specific structure via standard tools, which yields a hybrid system model that is well-suited to optimal control design. In particular, solving an OCP involving a hybrid system with PWA dynamics as an NLP has been proven to require shorter computation times and feature better scaling in the problem dimensions than standard approaches based on mixed-integer optimization [33, 28]. In addition, we recall that PWA regression is NP-hard in general [41], since it requires simultaneous classification of the $N$ samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ into modes, thereby calling for a regression of a submodel for each mode. By taking a NN-based perspective, we circumvent such a potentially challenging classification issue, thereby reducing the identification step to the training of the adopted NN architecture, a task that can be accomplished through standard gradient-based methods in view of the network’s output differentiability.
Remarkably, we obtain an LC model suitable for control as a direct consequence of a careful choice of a NN with specific architecture, thereby circumventing the requirement of strict complementarity to recover differentiability w.r.t. the main parameters as in [39, 16]. In contrast to [55], we give an explicit structure of a NN for which local stationarity conditions, coinciding with the standard KKT system, are known to hold for the MPCC obtained by embedding the NN as a hybrid model in an OCP.
The rest of the paper is organized as follows: we formalize the problem addressed in §2, and in §3 introduce our NN-based approach to system identification for control. In §4 we describe the OCP obtained when using our identified model structure, and develop our main theoretical results relating to that problem in §5. Finally, §6 reports a number of extensive numerical experiments.
@@ 2 Problem formulation
We will assume that we have available a finite collection of (state-input)-successor state measured triplets, $(x,u,x^{+})$, $x$, $x^{+}\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$ for an unknown but deterministic dynamic system. Our aim is to produce a data-driven model of this unknown system without running further experiments. Such a model shall be “descriptive enough”, i.e., it shall belong to a model class capable of capturing a wide variety of system behaviors, while at the same time being suited for optimal control.
To this end, we will consider throughout the paper the following finite horizon optimal control problem (OCP):
@@
where $x_{t}$, $t\in\mathcal{T}\coloneqq\{0,\ldots,T-1\}$, denotes the predicted system state $t$ timesteps into the future through the (possibly nonlinear) dynamical model $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, which is to be inferred from data and is characterized by parameters $\theta\in\mathbb{R}^{p}$. We denote by $\boldsymbol{u}\coloneqq\textrm{col}((u_{t})_{t\in\mathcal{T}})\in\mathbb{R}^{mT}$ the collection of control inputs to be chosen and by $\boldsymbol{x}\coloneqq\textrm{col}((x_{t})_{t\in\mathcal{T}\cup\{T\}})\in\mathbb{R}^{n(T+1)}$ the resulting state trajectory starting from the measured initial state $x_{0}=x(0)$. In addition, let $\ell_{t}$ and $\ell_{T}$ denote the stage and terminal costs respectively, which are assumed to meet the following standard condition:
@@
Thus, given $N$ samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, we aim at developing a system model $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that, once $\mathcal{N}_{\theta}$ is used within the OCP in (1), an optimal solution to the latter can be characterized via standard optimality conditions. To this end, we design $\mathcal{N}_{\theta}$ as a neural network (NN)-based representation of a hybrid system, ultimately leading to an OCP equivalent to (1) in the form of a mathematical program with complementarity constraints (MPCC) [44], whose Karush-Kuhn-Tucker (KKT) conditions [7, §5.1] will be necessary and sufficient to characterize an optimal solution.
For technical reasons, we will further assume that the state-input pair takes values in some bounded set:
@@
3 A neural network-based representation of hybrid systems
@@
To address the problem introduced in §2, we will design $\mathcal{N}_{\theta}$ as a two-layer NN with the simplified architecture illustrated in Fig. 1, and then make use of available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ to train the associated parameters, generically described by $\theta\in\mathbb{R}^{p}$. In particular, the underlying NN consists of an OptNet layer [2, 1], which takes the input pair $\textrm{col}(x,u)\eqqcolon y$ as a parameter to solve the following generic quadratic program (QP):
@@
thus returning an optimal solution $z^{\star}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{s}$ as output. This output is then passed through a linear layer with tailored matrix $W\in\mathbb{R}^{n\times s}$ and affine term $c\in\mathbb{R}^{n}$:
$x^{+}=Wz^{\star}(x,u)+c\eqqcolon\mathcal{N}_{\theta}(x,u).$ (3)
To ensure a unique optimizer for (2) so that there is no ambiguity in the definition of the dynamics in (3), during the training phase we will impose $Q\in\mathbb{S}^{s}_{\succ 0}$. This will be made possible in view of the main technical features possessed by the NN in Fig. 1, established next:
@@
The NN $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ in (3) enjoys the following properties:
@@
(ii) This part follows from Standing Assumption 2 and [32, Th. 2] directly. In fact, if one chooses $c=0$, and parameters $Q,q,A,b,G,h$ so that (2) becomes:
@@
with $Q\in\mathbb{S}^{s}_{\succ 0}$, then the structure in [32, Eq. 4] is immediately recovered. Specifically, the inequality constraints in (4) are obtained from those in (2) by simply imposing $G(y)=F$ and $h(y)=h-Gy$, for appropriate matrices $F$, $G$ and vector $h$ with row dimension $l\leq 2nr$ and $s\leq 2n$ columns (notice the slight abuse of notation). Remarkably, these choices are always possible since differentiability of the NN output w.r.t. $\theta$, which follows from the first property shown in this statement, implies that $\mathcal{N}_{\theta}$ is an end-to-end trainable NN. ∎
In the rest of the paper, when we will refer to the OptNet layer we will tacitly consider the structure in (4). The set of parameters characterizing $\mathcal{N}_{\theta}$ is then $\theta\coloneqq\{(Q,R,p,F,G,h),W,c\}\in\mathbb{R}^{p}$, which shall be determined during an offline training procedure using the available dataset $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$. With this regard, Proposition 1.(i) makes $\mathcal{N}_{\theta}$ in (3) an end-to-end trainable NN, meaning that one is actually able to set values for the parameters $\theta$ exploiting available samples, as commonly happens with any other NN. A differentiable output enables the use of a typical backpropagation strategy to find the gradient of the training loss function. Thus, one can train $\mathcal{N}_{\theta}$ with standard gradient descent methods for finding a minimum of that function.
As a main consequence of the second property in Proposition 1, instead, we note that (1) turns out to coincide with an OCP involving a continuous PWA dynamics. The latter is known to be equivalent to a number of hybrid system model classes [30], such as linear complementarity (LC) [39] or mixed-logical dynamical (MLD) models [6]. For this reason, the family of continuous PWA dynamics excels in capturing a wide variety of real-world system behaviours. In particular, since $\mathcal{N}_{\theta}$ will be trained on the basis of available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, (1) actually amounts to an OCP involving a data-based hybrid system representation.
Finally, note that the architecture expressed in (3)–(4) represents a totally valid NN in the machine learning realm, featuring attractive computational properties [8]. In particular, it incorporates an implicit layer [2, 8], which requires an iterative process to compute its output, unlike traditional explicit layers. Since the proposed method involves not only solving the QP in (4) but also learning its parameters, referring to it as a NN is consistent with the terminology used in the literature.